# Differentially Private Adaptation of Diffusion Models via Noisy Aggregated Embeddings

### v2

@@ C.2 Implementation
We follow the style guidance process introduced by Bansal et al. (2024), modifying it to include differential privacy mechanisms. Let $x_{c}$ denote the target style image, $x_{t}$ the noisy image at step $t$, and $\mathcal{E}(\cdot)$ the CLIP image encoder. The forward guidance process is defined as follows:
$\hat{\epsilon}_{\theta}(x_{t},y,t)=\epsilon_{\theta}(x_{t},y,t)+w\sqrt{1-\alpha_{t}}\nabla_{x_{t}}\ell_{\cos}(\mathcal{E}(x_{c}),\mathcal{E}(\hat{x}_{0})),$ (10)
@@
To integrate differential privacy, we encode each target image $x^{(i)}$ into $u^{(i)}=\mathcal{E}(x^{(i)})$ and aggregate these embeddings into $u^{*}$ using the centroid method. The aggregated $u^{*}$ guides the reverse diffusion process:
$\hat{\epsilon}_{\theta}(x_{t},y,t)=\epsilon_{\theta}(x_{t},y,t)+w\sqrt{1-\alpha_{t}}\nabla_{x_{t}}\ell_{\cos}(u^{*},\mathcal{E}(\hat{x}_{0})).$ (11)
@@
C.3 Ablation
@@
To better understand the limited effectiveness of style guidance in our experiments, despite its success in (Bansal et al., 2024), we applied our approach to a dataset of 143 paintings from Van Gogh’s Saint-Paul Asylum, Saint-Rémy collection(Innat, ) (Figure 9). Unlike the @eveismyname and Paris 2024 datasets, it is highly likely that Stable Diffusion has been trained on these images. Additionally,  Bansal et al. (2024) demonstrated successful adaptation towards the style of Van Gogh’s Starry Night as a single reference image, making this dataset a reasonable interpolation between their successful results and our more limited findings.
Appendix D Additional Style Transfer and Ablation Results
@@
Figure 11: Images generated by Stable Diffusion v1.5 using the prompt “A painting of Taylor Swift in the style of $<$@eveismyname$>$”, with the embedding $<$@eveismyname$>$ trained using DPAgg-TI (with different subsample sizes $m$) and TI with DP-SGD using different values of $\varepsilon$.
@@
Figure 12: Images generated by Stable Diffusion v1.5 using the prompt “An icon of the Eiffel Tower in the style of $<$Paris 2024 Pictograms$>$”, with the embedding $<$Paris 2024 Pictograms$>$ trained using DPAgg-TI (with different subsample sizes $m$) and TI with DP-SGD using different values of $\varepsilon$.
@@
Figure 13: Images generated by Stable Diffusion v1.5 using the prompt “An icon of a dragon in the style of $<$Paris 2024 Pictograms$>$”, with the embedding $<$Paris 2024 Pictograms$>$ trained using DPAgg-TI (with different subsample sizes $m$) and TI with DP-SGD using different values of $\varepsilon$.

### v3

@@ C.2 Implementation
We extend our approach to style guidance (SG) by leveraging the framework of Universal Guidance (Bansal et al. 2024). Specifically, we focus on CLIP-based style guidance, which optimizes the similarity between the CLIP embeddings of a target image and the generated image.
We encode each target image $x^{(i)}$ as $u^{(i)}$ via a CLIP image encoder, then aggregate the embeddings $u^{(1)},\dots,u^{(n)}$ into $u^{*}_{\text{DP}}$ using (5) or (7), depending on whether subsampling is applied. The aggregated embedding $u^{*}_{\text{DP}}$ is then incorporated into the reverse diffusion process as a style guide.
Let $x_{c}$ denote the target style image, $x_{t}$ the noisy image at step $t$, and $\mathcal{E}(\cdot)$ the CLIP image encoder. The forward guidance process is defined as follows:
$\hat{\epsilon}_{\theta}(x_{t},y,t)=\epsilon_{\theta}(x_{t},y,t)+w\sqrt{1-\alpha_{t}}\nabla_{x_{t}}\ell_{\cos}(\mathcal{E}(x_{t}),\mathcal{E}(\hat{x}_{0})),$ (10)
@@
To integrate differential privacy, we encode each target image $x^{(i)}$ into $u^{(i)}=\mathcal{E}(x^{(i)})$ and aggregate these embeddings into $u^{*}_{\text{DP}}$ using the centroid method. The aggregated $u^{*}_{\text{DP}}$ guides the reverse diffusion process:
$\hat{\epsilon}_{\theta}(x_{t},y,t)=\epsilon_{\theta}(x_{t},y,t)+w\sqrt{1-\alpha_{t}}\nabla_{x_{t}}\ell_{\cos}(u^{*}_{\text{DP}},\mathcal{E}(\hat{x}_{0})).$ (11)
@@
C.3 Style Transfer Results
C.4 Ablation
To better understand the limited effectiveness of style guidance in our experiments, despite its success in (Bansal et al. 2024), we applied our approach to a dataset of 143 paintings from Van Gogh’s Saint-Paul Asylum, Saint-Rémy collection(Innat 2022) (Figure 10). Unlike the @eveismyname and Paris 2024 datasets, it is highly likely that Stable Diffusion has been trained on these images. Additionally,  Bansal et al. (2024) demonstrated successful adaptation towards the style of Van Gogh’s Starry Night as a single reference image, making this dataset a reasonable interpolation between their successful results and our more limited findings.
@@ C.4 Ablation
Appendix D Copyright Protection
Modern generative models typically produce outputs via randomized sampling. Leveraging this inherent randomness, Vyas, Kakade, and Barak (2023) introduced Near Access-Freeness (NAF) as a metric to quantify the similarity between a model’s output and copyrighted content. The key idea is to compare the output distribution of a potentially infringing model to that of a safe model – one trained without access to the target content.
Formally, let $\mathsf{safe}$ be a mapping from a data point $x\in\mathcal{C}$ (where $\mathcal{C}$ is the collection of copyrighted samples) to a generative model $\mathsf{safe}(x)\in\mathcal{W}$ that is trained without using $x$. A canonical example is the leave-one-out-safe model, trained on the full dataset excluding $x$. Since $\mathsf{safe}(x)$ does not have access to $x$, the probability that it generates content resembling $x$ is exponentially small. Any such resemblance is considered fortuitous. Formally, the NAF criterion is defined as follows:
Definition 2 (Near Access-Freeness (Vyas, Kakade, and Barak 2023)).
Let $\mathcal{C}$ be a set of copyrighted samples and $\mathcal{W}$ a set of generative models. Given a mapping $\mathsf{safe}:\mathcal{C}\rightarrow\mathcal{W}$ and a divergence measure $\Delta$, we say a model $w$ is $k_{y}$-near access-free (or $k_{y}$-NAF) on prompt $y\in\mathsf{Y}$ if for every $x\in\mathcal{C}$,
$\Delta\left(p(\cdot|y)\,\|\,\mathsf{safe}_{x}(\cdot|y)\right)\leq k_{y}.$
If $k_{y}=0$, the model is indistinguishable from a safe model, meaning any resemblance to copyrighted material is by random chance. More generally, a small $k_{y}$ suggests the model is unlikely to generate outputs resembling $x$ with higher probability than a model that has never seen $x$.
D.1 Connection to Differential Privacy
NAF is closely related to concepts in Differential Privacy (DP) (Elkin-Koren et al. 2023). Depending on the divergence measure $\Delta$, NAF resembles different DP variants – for example, $\varepsilon$-DP when $\Delta=\Delta_{\mathrm{max}}$ (Dwork et al. 2006), and $(1,\varepsilon)$-Rényi DP when $\Delta=\Delta_{\mathrm{KL}}$.
Translating DP to generative models yields the following definition:
Definition 3 (Differentially Private Generation (DPG)).
Let $S$ and $S^{\prime}$ be neighboring datasets. Denote by $P_{S}(\cdot|y)$ the distribution over outputs generated by a model trained on, or adapted from, $S$ with algorithm $\mathcal{A}$, where randomness includes both training and generation stages. The generation is said to satisfy $\varepsilon$-Differentially Private Generation ($\varepsilon$-DPG) if for every $y\in\mathcal{Y}$,
$\Delta\left(P_{S}(\cdot|y)\,\|\,P_{S^{\prime}}(\cdot|y)\right)\leq\varepsilon.$
Here, neighboring datasets differ by a single data point (or privacy unit). If the training process is $\varepsilon$-DP, then the outputs naturally satisfy $\varepsilon$-DPG via the data processing inequality. One benefit of DPG is the flexibility to add noise during generation rather than training, potentially improving the utility-privacy tradeoff.
However, there are notable distinctions. $\varepsilon$-DP offers protection under arbitrary post-processing and multiple outputs, whereas $\varepsilon$-DPG only guarantees privacy for single outputs. Also, under DP, the trained model can be released, but under DPG, only the outputs are safe to share.
Elkin-Koren et al. (2023) highlight further differences: NAF is one-sided—comparing a model to a fixed safe reference—whereas DPG is symmetric. This asymmetry in NAF can enable better utility. Additionally, NAF allows more flexibility in choosing the safe model, which can be exploited in algorithm design.
Given these conceptual overlaps, both DP-SGD based training and our proposed private adaptation method DPAgg-TI satisfies $\varepsilon$-DP, so they naturally satisfy $\varepsilon$-NAF with the leave-one-out safe model.
We emphasize that this guarantee is meaningful only within the formal framework of NAF. It does not imply broader legal immunity or empirical indistinguishability from the original content. However, within this framework, satisfying $\varepsilon$-NAF allows us to argue that any close resemblance between outputs and private training data is no more likely than would be expected from a model that never had access to that data. This theoretical grounding supports the privacy and safety claims of our adaptation method.
Importantly, the goal of DPAgg-TI is to adapt to the style of a private image set—not its precise content. This distinction matters: style transfer is widely considered to fall under the doctrine of fair use, particularly in artistic and creative contexts. As discussed in Elkin-Koren et al. (2023) and further elaborated in legal analysis such as Carlini et al. (2023), generating new content in the style of a work, without reproducing its substantive elements, is generally not considered copyright infringement. Therefore, the use of DPAgg-TI to learn and reproduce stylistic attributes does not contradict the spirit or intent of the NAF framework. Instead, it offers a promising direction for responsibly fine-tuning generative models on private or copyrighted sources while respecting both privacy and intellectual property boundaries.
Appendix E Computational Cost Comparisons
Direct comparisons of computational cost across methods are inherently challenging due to differing training paradigms, optimization procedures, and parameter settings. Nonetheless, to provide a concrete sense of scale, we report representative computational costs for each method based on experiments conducted using a Stable Diffusion v1.5 model on a single NVIDIA A100 GPU. Below we summarize both training and inference overheads (the number of steps are optimized for each setup):
Method Steps Batch Size Time Memory Usage
TI (no DP) 10,000 (for 150 images) 1 25 min 7 GB
8 2.5 hours 20 GB
TI (DP-SGD) 30,000 (for 150 images) 1 80 min 7 GB
8 7 hours 20 GB
DPAgg-TI 2,000 per image N/A $\sim$5 min/image 7 GB
SG N/A N/A N/A N/A
Table 4: Training cost comparison across methods. Overhead from DP-SGD is relatively modest due to the low-dimensional embedding being optimized. N/A for SG means nothing is trained aside from the base model.
Method Steps Batch Size Time Memory Usage
TI (no DP, DP-SGD, DPAgg-TI) 50 1 1–2 sec 4 GB
100 1 1–2 min 58 GB
SG (no DP, DPAgg-SG) 500 1 $\sim$30 min 17 GB
Table 5: Inference cost comparison across methods.
Appendix F Additional Style Transfer and Ablation Results
@@
Figure 12: Images generated by Stable Diffusion v1.5 using the prompt “A painting of Taylor Swift in the style of <<@eveismyname>>”, with the embedding <<@eveismyname>> trained using DPAgg-TI (with different subsample sizes $m$) and TI with DP-SGD using different values of $\varepsilon$.
@@
Figure 13: Images generated by Stable Diffusion v1.5 using the prompt “An icon of the Eiffel Tower in the style of <<Paris 2024 Pictograms>>”, with the embedding <<Paris 2024 Pictograms>> trained using DPAgg-TI (with different subsample sizes $m$) and TI with DP-SGD using different values of $\varepsilon$.
@@
Figure 14: Images generated by Stable Diffusion v1.5 using the prompt “An icon of a dragon in the style of <<Paris 2024 Pictograms>>”, with the embedding <<Paris 2024 Pictograms>> trained using DPAgg-TI (with different subsample sizes $m$) and TI with DP-SGD using different values of $\varepsilon$.